OpenStudy (anonymous):

the expression (3^(1/3)/3^(-2/3)) is equivalent to (1) 1 (2) square root of three (3) 3 (4) (1/the cube root of three)

OpenStudy (anonymous):

it's 3. Here's why: (3^(1/3)/3^(-2/3)) = 3^(1/3) x [3^(-2/3)]^(-1) = 3^(1/3) x 3^(2/3) = 3^([1/3] + [2/3]) = 3^1 = 3.